OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. 6m=-36

OpenStudy (anonymous):

m= -6 what class are you in

OpenStudy (anonymous):

i do k12 online school and thanks

OpenStudy (kropot72):

Divide both sides by 6 and get: 6m/6=-36/6 m=-6